# The double gamma function and Vladimir Alekseevsky

### v2

@@ 2. Roots of functions similar to $\Gamma$ has a form
Visiting Leipzig in 1894, Alekseevsky published a short paper [Ale1894] in German in Leipz. Ber. (communicated by Sophus Lie), and, apparently, his work became known due to this publication. The paper is a sketch of [Ale1888]-[Ale1889] with some additional clams at the end of the paper. In particular,
@@
(1) [^8] These numbers are my.Functionen sind construirt aus einer Function $H$, die analog der Function $\Gamma$ ist.
@@
Clearly, in (3) Alekseevsky claims that an Appell function is a product of two multiple gammas [^9] In (1) ’Heine function’ means $q$-gamma, which can be expressed in $G(x;a)$ by the reflection formula (see our next subsection). Since we can express theta as product of 4 double gammas, we can express any double periodic function, see, e.g., [Akh1970], §19.
@@
In any case, the multiple gamma function with several parameters had not appear in published papers of Alekseevsky and actually was introduced by Barnes [Bar1904]. In Subsect. 3, we continue this discussion.
@@
$O(x):=\prod_{k=1}^{\infty}(1-q^{2k}e^{2\pi ix}),\qquad\text{where $q=\pi i\alpha$,}$
(this follows from the reasoning of the previous subsection). Applying the Jacobi triple identity (see, e.g., [Akh1970], S 18), we can easily express the standard theta functions in the double gamma.
1.7. A counterpart of the $\psi$-function. Let us return to the work [Ale1888]-[Ale1889]. Recall that the classical $\psi$-function (see, e.g., [Erd1953], 1.7) is defined by
[…]
@@
for $a=1$ this gives (1.6). So, here he has not come to a new special function, but he found some interesting formulas.
@@ 2. On the history of the double gamma function
2.1. The work by Kinkelin, 1860. Hermann Kinkelin [Kin1860] introduced a function [^10] His original notation is ${\bf G}(x)$. $K(x)$ satisfying the identity
@@
Kinkelin obtained some properties of the function $K(x)$. In particular, he wrote its representations as limits of products and expression of elementary integrals in terms of $K(x)$ (§§10-12). The most remarkable of his identities for $K(x)$ is the multiplication formula, which was established in a better form than was done later by Alekseevsky:
@@
So, the constant $\widetilde{\omega}$ is connected with $G\bigl{(}\frac{1}{2}\bigr{)}$, which appears in many formulas for the double gamma, by a simple relation. Passing to the limit as $n\to\infty$ in (2.3), we get
[…]
Also, he expressed $\widetilde{\omega}$ in terms of the zeta-function $\zeta(x)=\sum_{n=1}^{\infty}n^{-x}$,
@@
where $\gamma$ is the Euler constant (see §8). Another expression is given by the following counterpart (§9) of Raabe formula (see [Erd1953], 1.9(19)) for the gamma function:
[…]
(this formula was later repeated by Alekseevsky, §10).
@@
The expression $A:=(\widetilde{\omega})^{1/2}e^{1/12}$ now is called the Glaisher–Kinkelin constant, cf. Glaisher [Gla1877].
The work by Kinkelin didn’t attract immediate attention and became a subject of interest only in the second half of the 1890s. The problem (2.1) seems even more strange than (1.1), so it is natural to ask about his way to the ’Kinkelin function’.
@@
$\sum_{j=0}^{n-1}B_{p}\Bigl{(}x+\frac{j}{n}\Bigr{)}=j^{1-p}B_{p}(nx)$
@@
Notice that now $f_{q}(x)$ satisfies the difference equations $f_{q}(x+1)-f_{q}(x)=\frac{1}{x^{q+1}}$. This family also was not ‘new’.
However, he tries to vary the equality (2.4), the function $f(x)=\ln\Gamma(x)$ satisfies the addition theorem
@@
which is equivalent to the multiplication formula for $\Gamma$-function. Also, it satisfies the difference equation
@@
[…]
@@
Correcting $F(x)\to\widetilde{F}(x)$ by an appropriate quadratic summand, we can reduce the last equation to the form
@@
2.2. The double gamma function, 1899-1907. The work of Alekseevsky became known (may be, due to the publication [Ale1894]). After this, the double gamma function was a topic of publications by Ernest Barnes [Bar1899], [Bar1900.1], [Bar1900.2], [Bar1901], [Bar1904], Jean Beaupain [Bea1903], [Bea1904], [Bea1907], Godfrey Hardy [Har1905.1], [Har1905.2], Vladimir Steklov [Ste1904]. The most important result of these works was the introduction of the multiple gamma functions by Barnes [Bar1904].
@@
Alekseevsky [Ale1904] showed that systems of functions $G_{n}(x)$ (see above Subsect. 1) and $K_{n}(x)$ can be expressed one from another,
@@
As it was observed in [Ada2014], this function was mentioned in the second edition [WhWa1912] of E. T. Whittaker, G. N. Watson A course of Modern Analysis (Chapter XII, Examples 48-50, namely (1.6) as a definition, (1.1), (1.5), (1.11)). Next, it was mentioned in the fourth edition [Grad1963] of Gradshteyn, Ryzhik Tables of integrals, sums, and series, lines 6.441, 8.333, with references to Whittaker–Watson.
The double gamma is absent in other handbooks on special functions as Higher transcendental functions [Erd1953], or Prudnikov, Brychkov, Marichev, [PBM1985].
I have seen (mainly, due to [Gus1952]) few additional appearances of such functions in literature. L. Bendersky [Ben1933] published in 1933 big detailed investigation on higher Kinkelin functions (he probably didn’t know about his predecessors). Boris Rimsky-Korsakov published the paper [Rim1947] on multiplication theorems for higher Kinkelin functions [^11] The work is known to me only through references in [Gus1950], [Gus1952]. I don’t present this unusual formula, since it contains typos, which I am unable to correct., and the paper [Rim1957] about further generalizations of the gamma function.
The double gamma function is discussed in the obituary of Barnes written by Edmund Whittaker [Whi1954].
2.5. A new appearance. The double gamma function came again in the work by Takuro Shintani [Shi1976]-[Shi1977]. After that, the function started showing up in various math problems, we only present few references [Vig1979], [Sar1987], [Var1988], [NiYo2009], [Ada2014], [Ost2013], [AlKu2023], [KarKu2024].
@@
(a special case of the double sine corresponding to equal periods was considered by Otto Hölder [Hol1886]). This function is important by itself. Much later (and independently) Ludvig Faddeev [Fad1995] introduced ’quantum dilogarithm, which is very similar to the double sine. However, it appeared in a completely new context, and now the term ’quantum dilogarithm’ is even more common than ’double sine’.
@@
For instance, see nice applications of the double sine in representation-theoretical spirit (which were the reason of my interest to this topic) in [Kash1998], [FoCh1999], [KLS2002]).
@@
2.6. Why was the double gamma forgotten 70 years? I’ll try to speculate on possible causes. One of the reasons is clear, the function was seriously different from special functions, which were considered at that time, and had no immediate applications. In my opinion, there were some accidental reasons.
Firstly, both Alekseevsky (see below) and Barnes (1874-1953), see [Whi1954], left mathematics.
Secondly, this can be partially related to the structure of [Ale1888]-[Ale1889]. The work can seem like a free fantasy, real external relations become clear only at the end of the second paper. Also, $q$-hypergeometric functions were almost unknown in Russia until the appearance for interest to representations of classical groups over finite and $p$-adic fields in the 1960s.
@@
Thirdly, a meticulous reader can decide that the actual topic of the paper are the functions $\ln G(x)$, $\ln G_{n}(x)$, $\ln G(x;a)$. For instance, in the paper, the double gamma itself (without a logarithm) never appears in the right hand sides of integrals, this contrasts with the behavior of $\Gamma(x)$) (the gamma function is usual in right hand sides of definite integrals, see, e.g. [Grad1963], [PBM1985]; on the other hand an appearance of $\ln\Gamma(\cdot)$ is relatively exotic). Integrals as (1.10), (1.11) are not too impressive and admit other transformations. Only in §22 a reader can understand that products of double gammas can participate in Heine’s theory of $q$-hypergeometric functions.
@@
Now there are no reasons for such doubts. Ludvig Faddeev [Fad1995] observed that the double sine is a kernel of a natural integral operator in $L^{2}(\mathbb{R})$ satisfying highly nontrivial identities, see [Fock1997], [Kash1998], [FoCh1999]. See also the work [KLS2002], where a certain spectral measure for quantum groups are expressed as products of the double sines. Recall that products of usual gamma functions appear as spectral densities for differential and difference operators (as far as they admit explicit evaluations), also such products appear as spectral densities for unitary representations of semisimple groups.
@@
In several papers 1899-1907, there was used the term ’Alexeiewsky function’. The widely used modern term ’Barnes double gamma’ arises to Shintani [Shi1976][Shi1977], who found old Barnes works. Alekseevsky, who was mainly the author of one big work [Ale1888], [Ale1889], [Ale1894], [Ale1904], at that time was forgotten, Barnes was well known due to ’Barnes integrals’, which are a fundamental subject and a tool in theory of special functions, see, e.g., [Whi1954], [AAR1999]. Notice also that Barnes type integrals and the convolution formula for the Mellin transform were the main tools for production of tables of integrals of A. P. Prudnikov, Yu. A. Brychkov, O. I. Marichev [PBM1985] (see explanation of the technology in [Mari1978]).
3. Biography of Alekseevsky
As we have seen above, the main work by Alekseevsky was forgotten, he also was forgotten as a mathematician (his name was mentioned in a few Russian special researches on the history of mathematics, [Gus1950], [Gus1952], [Kru1967], [Bakh1967], [Kru1967.1], [Bel2000] [^12] I have not seen the last book.). On the other hand, he left his memory as an effective director (1907-1911) of the Tomsk Technological Institute.
3.1. Historiography. The main sources about Alekseevsky known for me are:
@@ 3. Biography of Alekseevsky
— Obituary written by president (1906-1946) of the Kharkov mathematical society Dmitry Sintsov [Sin1917].
@@
— Review by Steklov and Sintsov [StSi1904] of works of Alekseevsky (apparently written by Steklov) supporting application of Alekseevsky to the position of extra-ordinary professor.
@@
— Dissertation by Victor Gussov [Gus1950] on the history of investigations on the gamma function and Bessel functions [^13] Aleksey Markushevich was an opponent of the thesis and used it in his essay on history of complex analysis [Mark1951]. However, he had not mentioned Alekseevsky in ’Essays’, and in his treatise on functions of complex variable. and his paper [Gus1952].
— Works based on archive documents about the history of science in Tomsk and on the history of Tomsk Technological Institute, mainly, Nikolay Krulikovsky [Kru1967] and Yulia Granina [Gran2003]. The latter paper contains a lot of biographical information, below we do not try to abstracting it.
3.2. Early biography. Vladimir Petrovich Alekseevsky was born 07(19).04.1858 in Novgorod governorate (‘guberniya’) and baptized in the village Peretënka (near the railway station Okulovka on the line Moscow–St.Petersburg). His father, Petr Osipovich Alekseevsky, was ‘praporshchik’ [$\simeq$ warrant officer] of the Corps of transport engineers [Корпус инженеров путей сообщения]. In 1877, Alekseevsky graduated from the gymnasium in Nizhny Novgorod [^14] Aleksandr Lyapunov, Ivan Privalov, physicist Grigory Landsberg, zoologist Aleksandr Formozov, philosopher Semyon Frank, chemist Alexey Favorsky graduated from the same gymnasium in different years..
3.3. The Kharkov University, 1877-1905. After gymnasium, Alekseevsky entered the Natural Science Department of the Faculty of Physics and Mathematics of the Kharkov University. Two years later, he transferred to Mathematical Department. In 1881, ‘for family reasons’, he left the University [^15] The main obstacle to his research activities was his difficult financial situation’ as was told in [Bakh1967] (and this is clear from [Sin1917]). However, Sintsov in the obituary does not indicate any details (and they are absent in other sources).. In 1884, he passed examinations without attending lectures and graduated from the university. At that moment he had three published papers [Ale1884.1]-[Ale1884.3] on integrability of certain ordinary differential equations. They were counted as a graduate (‘candidate’) work [^16] There is no information about his scientific adviser(s). In [Ale1884.1]-[Ale1884.3], he mentioned the names of Vasily Imshenetsky, who worked at the Kharkov University 1877-1881 (and moved to St.-Petersburg after this), and Aleksey Letnikov, who worked in Moscow..
Since 1881, Alekseevsky taught mathematics, mechanics, physics in Kharkov in the real school (until 1891) and in the Second women’s gymnasium (until 1887). In 1887, he received a scholarship to prepare for a professorship (an analog of a postgraduate student). His main work was published in [Ale1888][Ale1889] (in fact, it is one paper split in the middle of a sentence). In 1892, this work was presented as a dissertation171717 A ‘magister dissertation’ in the terminology of that time. Apparently, in modern terminology, it corresponds to a PhD. [Ale1892]. After the defence (21.02.1893), he became a privatdozent in the Kharkov University.
In March 1894 – November 1895 Alekseevsky visited Leipzig, Göttingen, and Paris. At that time, he published the paper [Ale1894] in Leipz. Ber..
After coming back, he decided to leave the university and became a teacher in a gymnasium in the town Starobelsk in Kharkov governorate since ’university …could do nothing to support the scientist it highly estimated’ (a quotation of a rapport of Steklov in [Sin1917], who worked at the Kharkov university in that time). Alekseevsky returned to the university the next year. Starting 1898, he also worked at the Kharkov Technological Institute.
Four pages of the obituary [Sin1917] (and big part of [StSi1904], see, also, the correspondence of Steklov and Lyapunov in [Ste1991]) contain a description of efforts 1896-1904 of several colleagues to transfer Alekseevsky to the position of extraordinary professor. Participants were: Aleksandr Lyapunov, Vladimir Steklov, Matvey Tikhomandritsky (specialist in elliptic functions), Konstantin Andreev, Dmitry Grave, Ludwig Struve (astronomer), Ivan Osipov (chemist). The nature of obstacles is not clear from the published literature. Steklov (letter to Lyapunov, 14.11.1903) characterized this as an ’absurd story’.
From [Sin1917]: Steklov [1902] explains, why V. P.’s doctoral dissertation [$\simeq$ habilitation] was delayed – ’he could have compiled from the scientific results a fully satisfactory dissertation’ and expresses his conviction that ‘if V. P. did not take care to acquire formal rights, having at his disposal all the scientific materials, this circumstance can hardly diminish the scientific merits of V. P. Alekseevsky’. ‘But we must say even more. By now, V. P. Alekseevsky has prepared for printing and is already printing an extensive course of differential and integral calculus.…V. P. Alekseevsky’s course is not a template course adapted to the usual programs of higher educational institutions. It is an independent academic work, where, along with pedagogical requirements, it combines a rigorous and modern approach to the study of the subject’ (finally, this book was not published).
After fifth attempt, 12.01.1904 Alekseevsky got the position of acting extraordinary professor in Kharkov. In the next year, he moved to Tomsk for full professorship.
@@
3.4. The multiple gamma function and a lost work. Let us return to the discussion of Subsect. 1. In the review of Alekseevsky works [StSi1904], Steklov and Sintsov wrote:
@@
And at present time, Alekseevsky continues to successfully develop his theory of functions, which he now calls gammamorphic. He came to functions of an even more general type, for which Kinkelin functions [Steklov keeps in mind $G_{n}(x)$] are limit cases, and which are related to the known Riemann $\zeta(s)$ and many other functions important in Analysis and its applications.
@@
These recent studies were reported at the meetings of the Kharkov mathematical society in 1902 and 1903. At the present time, part of them is printing in Communications of our society.
@@
We see in [Khar1904] the following titles of talks by Alekseevsky on the Kharkov mathematical society:

### v3

@@ 2. Roots of functions similar to $\Gamma$ have the form
Visiting Leipzig in 1894, Alekseevsky published a short paper [Ale1894] in German in Leipz. Ber. (communicated by Sophus Lie), and his work apparently became known due to this publication. The paper is a sketch of [Ale1888]-[Ale1889] with some additional claims at the end of the paper. In particular,
@@
(1) [^18] These numbers are my addition.Functionen sind construirt aus einer Function $H$, die analog der Function $\Gamma$ ist.
@@
Clearly, in (3) Alekseevsky claims that an Appell function is a product of two multiple gammas [^19] In (1) ’Heine function’ is (3.20); it is the $q$-gamma function up a normalization, which can be expressed in $G(x;a)$ by the reflection formula (see my next subsection). Since we can express theta as a product of 4 double gammas, we can express any double periodic function, see for example [Akh1970], §19.
@@
In any case, the multiple gamma function with several parameters had not appear in Alekseevsky’s published papers and actually was introduced by Barnes [Bar1904]. In Subsect. 5, we continue this discussion.
@@
(3.20) $O(x):=\prod_{k=1}^{\infty}(1-q^{2k}e^{2\pi ix}),\qquad\text{where $q=\pi i\alpha$,}$
(this follows from the previous subsection’s reasonings). Applying the Jacobi triple identity (see, for example [Akh1970], §18), we can easily express the standard theta functions in the double gamma.
3.7. A counterpart of the $\psi$-function. Let us return to the work [Ale1888]-[Ale1889]. Recall that the classical $\psi$-function (see, for example [Erd1953], 1.7) is defined by
$\psi(z)=\frac{d}{dz}\ln\Gamma(z)=\frac{\Gamma^{\prime}(z)}{\Gamma(z)}=\int_{0}^{\infty}\Bigl(\frac{e^{-t}}{t}-\frac{e^{-zt}}{1-e^{-t}}\Bigr)\,dz.$
@@
for $a=1$ this gives (3.6). So, here he has not come to a new special function, but he has found some interesting formulas.
@@ 4. On the history of the double gamma function
4.1. Kinkelin’s work, 1860. Hermann Kinkelin [Kin1860] introduced a function [^20] His original notation is ${\bf G}(x)$. $K(x)$ satisfying the identity
@@
Kinkelin obtained some properties of the function $K(x)$. In particular, he wrote its representations as limits of products and expressions of elementary integrals in terms of $K(x)$ (§§10-12). The most remarkable of his identities for $K(x)$ is the multiplication formula, which was established in a better form than Alekseevsky’s (3.8):
@@
So, the constant $\widetilde{\omega}$ is connected to $G\bigl(\frac{1}{2}\bigr)$, which appears in many formulas for the double gamma, by a simple relation. Passing to the limit as $n\to\infty$ in (4.3), we get
$\ln\widetilde{\omega}:=2\int_{0}^{1}\ln K(x)\,dx.$
He also expressed $\widetilde{\omega}$ in terms of the zeta-function $\zeta(x)=\sum_{n=1}^{\infty}n^{-x}$,
@@
where $\gamma$ is the Euler constant (see §8). Another expression is given by the following counterpart (§9) of the Raabe formula (see [Erd1953], 1.9(19)) for the gamma function:
$\int_{x}^{x+1}K(t)\,dt=\tfrac{1}{2}\ln\widetilde{\omega}+\tfrac{1}{4}x^{2}(2\ln x-1)$
(Alekseevsky later repeated this formula).
@@
The expression $A:=(\widetilde{\omega})^{1/2}e^{1/12}$ is now called the Glaisher–Kinkelin constant, cf. Glaisher [Gla1877].
Kinkelin’s work didn’t attract immediate attention and became a subject of interest only in the second half of the 1890s. The problem (4.1) seems even more artificial than Alekseevsky’s (3.1), so it is natural to ask about his way to the ’Kinkelin function’. Apparently, it was a result of the following ’speleology’.
@@
$\sum_{j=0}^{n-1}B_{p}\Bigl(x+\frac{j}{n}\Bigr)=n^{1-p}B_{p}(nx)$
@@
Notice that now $f_{q}(x)$ satisfies the difference equations $f_{q}(x+1)-f_{q}(x)=\frac{1}{x^{q+1}}$. This family also was not ‘new’ in Kinkelin’s time.
However, he tries to vary the condition (4.4). The function $f(x)=\ln\Gamma(x)$ satisfies the addition theorem
@@
which is equivalent to the multiplication formula for the gamma function. Also, it satisfies the difference equation
@@
We take primitives from both sides of identities (4.6), (4.7) and get two equations for the primitive $F(x):=\int f(x)\,dx$. Equation (4.7) yields
@@
Correcting $F(x)\mapsto\widetilde{F}(x)$ by an appropriate quadratic summand, we can reduce the last equation to the form
@@
4.2. The double gamma function, 1899-1907. The work of Alekseevsky became known (perhaps due to the publication [Ale1894]). After this, the double gamma function was a topic of publications by Ernest Barnes [Bar1899], [Bar1900.1], [Bar1900.2], [Bar1901], [Bar1903], [Bar1904], Jean Beaupain [Bea1903], [Bea1904], [Bea1907], Godfrey Hardy [Har1905.1], [Har1905.2], and Vladimir Steklov [Ste1904]. The most important result of these works was Barnes’s introduction of the multiple gamma functions [Bar1904].
In particular, the third Barnes’s publication [Bar1899] and his further works [Bar1900.1][Bar1903] were devoted to the double gamma. In [Bar1903] he found the first application: he expressed the electrostatic capacity coefficients of a pair of non-intersecting spheres in terms of the double gamma.
In the work [Bar1904] he introduced a new topic, namely the multiple gamma function; it is a function $G_{n}(x)$ depending on parameters $\omega_{1}$, …, $\omega_{n}$ and satisfying functional equations
where $A_{n}$ are certain elementary expressions and $G_{n-1}$ is the multiple gamma of the previous level.
Barnes became famous due to his later works of 1907–08 on the ’Mellin–Barnes integrals’, i.e., integrals of the type
$\int_{-i\infty}^{\infty}z^{-s}\frac{\prod\Gamma(a_{k}+s)\prod\Gamma(b_{l}-s)}{\prod\Gamma(c_{m}+s)\prod\Gamma(d_{n}-s)}\,ds.$
Integrals of such kind appeared earlier in works by Salvatore Pincherle and Hjalmar Mellin. In any case, Barnes’s research was a breakthrough — thanks to him such integrals became a powerful and flexible tool of theory of special functions, see for example [Whi1954], [AAR1999], [Mari1978] (in particular, Mellin–Barnes integrals were the main tool of creations of tables of integrals [PBM1985], see [PBM1989]; see also the impressive table of Mellin-Barnes integrals in volume 3 of [PBM1985], Chapter 8).
@@
Alekseevsky [Ale1904] showed that systems of functions $G_{n}(x)$ (see above Subsect. 3) and $K_{n}(x)$ can be expressed through each other,
@@
As was observed in [Ada2014], this function was mentioned in the second edition [WhWa1912] of E. T. Whittaker, G. N. Watson A course of Modern Analysis (Chapter XII, Examples 48-50, namely (3.6) as a definition, (3.1), (3.5), (3.11)). Next, it was mentioned in the fourth edition [Grad1963] of Gradshteyn, Ryzhik Tables of integrals, sums, and series, lines 6.441, 8.333, with references to Whittaker–Watson.
The double gamma was absent in other handbooks on special functions as Higher transcendental functions [Erd1953], or Prudnikov, Brychkov, Marichev, [PBM1985].
I have seen (mainly, due to [Gus1952]) a few other appearances of such functions in literature. In 1933, L. Bendersky [Ben1933] published a large detailed investigation on higher Kinkelin functions (he probably didn’t know about his predecessors).
Boris Rimsky-Korsakov published a paper [Rim1947] on the reciprocity theorem for higher Kinkelin functions:
$\frac{\Bigl\{\prod\limits_{k=0}^{q-1}K_{n}\bigl(pz+\frac{kp}{q}\bigr)\Bigr\}^{q^{n}}}{\Bigl\{\prod\limits_{k=0}^{p-1}K_{n}\bigl(qz+\frac{kq}{p}\bigr)\Bigr\}^{p^{n}}}=\Bigl(\frac{p}{q}\Bigr)^{\frac{p^{n}}{n+1}\sum_{k=0}^{n-1}B_{n+1}\bigl(qz+\frac{kq}{p}\bigr)}\pi_{n}^{(q^{n+1}-p^{n+1})/2},$
where $B_{m}(\cdot)$ are the Bernulli polynomials and $\pi_{n}$ are certain constants. For $n=1$ this statement is the Schobloch reciprocity formula ([Scho1884], [Srin2007]) for the gamma function. He also published the paper [Rim1957] and PhD thesis [^21] I have not seen it. [Rim1957.1] about further generalizations of the gamma function.
The double gamma function was discussed in the obituary of Barnes, written by Edmund Whittaker [Whi1954].
4.5. A new appearance. The double gamma function appeared again in the work of Takuro Shintani [Shi1976]-[Shi1977]. After that, the function started showing up in various math problems, I only present a few references [Vig1979], [Sar1987], [Var1988], [NiYo2009], [Ada2014], [Ost2013], [AlKu2023], [KarKu2024].
@@
(a special case of the double sine corresponding to equal periods was considered by Otto Hölder [Hol1886]). This function is important by itself. Much later (and independently) Ludvig Faddeev [Fad1995] introduced the quantum dilogarithm, which is very similar to the double sine. However, it appeared in a completely new context, and the term ’quantum dilogarithm’ is now even more common than ’double sine’.
@@
For instance, see some nice applications of the double sine in representation-theoretical spirit (which were the reason of my interest in this topic) in [Kash1998], [FoCh1999], [KLS2002]).
@@
4.6. Why were the double gamma and the multiple gamma forgotten 70 years? I’ll try to speculate on possible causes. One of the reasons is clear: the function was different significantly from special functions that were considered at that time. In my opinion, there were also some accidental reasons.
Firstly, both Alekseevsky and Barnes (1874-1953) left mathematics. Alekseevsky became a professor of mechanics and a rector of a technological institute (his last research publication was in 1904). Barnes became an Anglicanic deacon, a theologist, and later a bishop of Birmingham, two of his last research publications are dated 1908 and 1910, see [Whi1954]).
Secondly, this can be partially related to the structure of [Ale1888]-[Ale1889]. The work can seem like a free fantasy, since its real external relations become clear only at the end of the second paper. There was a chance for a new appearance of the double gamma after a historical-mathematical paper by Viktor Gussov [Gus1952], which contained a good exposition of [Ale1888]-[Ale1889] (I think many Russian mathematicians have seen this article and were surprised reading it). But that didn’t happen.
@@
Thirdly, a meticulous reader could decide that the actual topic of the works on the double gamma are the functions $\ln G(x)$, $\ln G_{n}(x)$, $\ln G(x;a)$. For instance, in Alekseevsky’s papers, the double gamma itself (without a logarithm) never appears in the right hand sides of integrals, which contrasts with the behavior of $\Gamma(x)$. The classical gamma function appears in the right hand sides of a huge number of definite integrals, see for example [Grad1963], [PBM1985]; on the other hand, appearances of $\ln\Gamma(\cdot)$ in the right hand sides of integrals are relatively rare. Next, integrals like (3.10), (3.11) are not too impressive (the right hand side is not better than the left hand side) and admit other transformations. Only in §22 of [Ale1888][Ale1889] a reader can understand that products of the double gammas can participate in Heine’s theory of $q$-hypergeometric functions [^22] Also, $q$-hypergeometric functions were almost unknown in Russia until the appearance of interest in representations of classical groups over finite and $p$-adic fields in the 1960-s. So, to a Russian reader, motivation of [Ale1888]-[Ale1889] was unclear.,
@@
Now there are no reasons for such doubts. Ludvig Faddeev [Fad1995] observed that the double sine is a kernel of a natural integral operator in $L^{2}(\mathbb{R})$ satisfying highly nontrivial identities, see [Fock1997], [Kash1998], [FoCh1999]. See also the work [KLS2002], where certain spectral measures for quantum groups are expressed as products of the double sines. Recall that products of the usual gamma functions appear as spectral densities for differential and difference operators (as far as they admit explicit evaluations), such products also appear as spectral densities in numerous problems of harmonic analysis for semisimple Lie groups.
@@
The term ’Alexeiewsky function’ was used in several papers published in 1899-1907. The widely used modern term ’Barnes double gamma’ was originated by Shintani [Shi1976][Shi1977], who found old Barnes works. Alekseevsky, who was mainly the author of one big work [Ale1888], [Ale1889], [Ale1894], [Ale1904], was forgotten at that time, while Barnes was well known due to the Barnes integrals (see above).
5. Alekseevsky: his historiography, lost work, and publications
5.1. Historial sources and historiography. I used the following contemporary historical sources on the biography of Alekseevsky:
@@ 5. Alekseevsky: his historiography, lost work, and publications
— Obituary written by the president (1906-1946) of the Kharkov Mathematical Society Dmitry Sintsov [Sin1917].
@@
— Steklov and Sintsov’s review [StSi1904] of Alekseevsky’s works (apparently written by Steklov) supporting Alekseevsky’s application to the position of extra-ordinary professor in Kharkov University.
— Steklov’s correspondence published in [Ste1991].
As told above, Alekseevsky’s work was forgotten. But his name several times was mentioned in histories of mathematics and education in Russia.
@@
— Viktor Gussov wrote an interesting and thorough dissertation [Gus1950] on investigations of the gamma function and the Bessel functions in Russia (the work was published in [Gus1952][Gus1953]). He found Alekseevsky’s works and exposed them in [Gus1950], [Gus1952]. Aleksey Markushevich was an opponent of Gussov’s thesis and used it in his ’Essays’ on the history of complex analysis [Mark1951]. However, he did not mention Alekseevsky in his ’Essays’, nor in his historical investigation [Mark1981], nor in his famous treatise Theory of functions of complex variable. Clearly, Gussov’s paper [Gus1952] was used in the publications mentioned below.
— Some information is contained in Èsfir Bakhmutskaya’s work [Bakh1967] on mathematics in Kharkov University.
— The Tomsk period of Alekseevsky’s biography was well-investigated in several works about the history of science in Tomsk (see Nikolay Krulikovsky [Kru1967], [Kru1967.1], Yulia Granina [Gran2003], Belomestnyh [Bel2000]).
@@
5.2. The multiple gamma function and a lost work. Let us return to the discussion in Subsect. 3. In their review of Alekseevsky works [StSi1904], Steklov and Sintsov wrote:
@@
Now as well, Alekseevsky continues to successfully develop his theory of functions, which he now calls gammamorphic. He came to functions of an even more general type, for which Kinkelin functions [Steklov keeps in mind $G_{n}(x)$] are limit cases, and which are related to the known Riemann $\zeta(s)$ and many other functions important in Analysis and its applications.
@@
These recent studies were reported at the meetings of the Kharkov Mathematical Society in 1902 and 1903. Part of them are currently beeing printing in our society[s Communications.
@@
We see in [Khar1904] the following titles of Alekseevsky’s talks at the Kharkov Mathematical Society: